Terry is mixing a punch which contains 3 liters of juice for every 2 liters of soda. If he wants to make 21
liters of punch, how much juice does he need

Answer by lwsshak3(11628)

Terry is mixing a punch which contains 3 liters of juice for every 2 liters of soda. If he wants to make 21
liters of punch, how much juice does he need
***
let x=amt of juice needed
21-x=amt of soda in punch
..
x/(21-x)=3/2
63-3x=2x
5x=63
x=63/5
amt of juice needed=63/5=12.6 liters